OpenStudy (anonymous):

How many solutions are there to the following system of equations? 2x – y = 2 –x + 5y = 3

OpenStudy (anonymous):

Just 1 solution x=13/9 y=8/9 Since they are linear and intersect at one location

OpenStudy (anonymous):

ok thnx :)